# Advanced Encryption Standard AI artificial intelligence

### v2

@@ 4.5. Guided Symbolic Execution
Starting at the cycle of interest, our guided SE approach first constructs the Boolean function describing the combinational sub-graph in front of a net. Next, it replaces all previously identified control signals with concrete values from simulation while keeping all other variables symbolic. Each variable left in the Boolean function now corresponds to an output net of a sequential gate. If the sequential gate is not one of our endpoints, it finds the last cycle in which the gate was updated and continues back-tracing from the data inputs of the gate at that cycle. This process is repeated until the resulting Boolean function only contains variables corresponding to the output nets of our endpoints. To resolve potential recursive dependencies, our guided SE approach introduces intermediate variables to break up these dependencies when they are detected. To achieve this, we automatically identify sequential loops within the netlist and assign intermediate variables to the registers at the outputs of these loops.
@@
Applying guided SE is highly specific to the target implementation and the goals of the reverse engineer. Therefore, it is impossible to express its effectiveness in numbers. Still, we applied guided SE in a white-box case study and verified the results against a ground truth in Section 4.6 to demonstrate correctness. In a second evaluation step, we applied guided SE on the Maggie netlist to highlight its efficiency and scalability. To this end, we semi-automatically extracted a Python script describing the circuit’s algorithmic behavior during execution, after initially performing the same task manually in Section 3.3. Thereby, we reduced the required effort from multiple weeks to merely a matter of days.
@@ 4.6. White-Box Case Study
To assess the applicability of our workflow, we conducted a white-box case study on a signal-processing design similar to Maggie. This additional study is conducted because it enables a comparison between the recovered design against a known ground truth, which is impossible in the black box setting. To this end, we synthesized an open-source Hilbert transformer design (OpenCores, [n. d.]) for a Xilinx 7-series FPGA and then extracted an algorithmic abstraction from the resulting gate-level netlist using only our generalized techniques.
@@
The initial netlist comprised 425 combinational and 619 sequential gates. After netlist pre-processing, we ended up with 1025 gates in total, 374 of them combinational and 651 sequential. Next, we recovered word-level structures using our generalized algorithms. The algorithm classified all 17 carry chains into 10 subtractions and 7 additions, identified 31 16-bit registers, and annotated 135 multi-bit pin groups to registers and arithmetic structures. When comparing against the ground truth, DANA achieved an NMI of 0.99, and our bit-order propagation assigned a correct bitorder to all pin groups. In total, we automatically assigned over 90% of all gates to word-level structures.
@@
ibex Lattice 5 0 3 0 0 0 1 9 0.04 0.96 0.97 0.97 0.97 165 0.25 0.39 0.91
Xilinx 3 0 0 0 0 0 3 6 0.01 0.98 0.98 0.96 0.97 148 0.16 0.86 1.00
icicle Xilinx 1 1 4 0 0 0 1 7 0.05 0.91 0.95 0.94 0.95 345 0.79 0.90 0.84
Lattice 3 0 6 0 0 1 0 10 0.12 0.86 0.94 0.90 0.94 112 0.31 0.62 0.94
simple_risc_v Xilinx 3 1 4 0 0 2 0 10 0.07 0.97 0.99 0.98 0.98 117 0.29 0.91 1.00
Lattice 2 1 4 0 0 6 1 14 0.13 0.98 0.99 0.98 0.97 136 0.32 0.95 1.00
canny_edge_detector Xilinx 57 0 1 2 49 3 6 118 0.77 0.81 0.81 0.61 0.64 619 0.46 0.91 0.99
Lattice 69 0 1 0 34 6 8 118 0.60 0.81 0.85 0.71 0.82 813 0.43 0.78 1.00
fft64 Xilinx 25 5 0 0 4 0 6 40 0.27 0.90 0.97 0.79 0.96 351 0.37 0.97 0.97
Lattice 93 4 5 0 6 0 1 109 0.40 0.88 0.94 0.77 0.91 604 0.59 0.93 0.96
sha256 Xilinx 10 0 0 0 0 0 3 13 0.26 0.91 0.96 0.75 0.87 134 0.25 0.97 1.00
Lattice 11 0 1 0 0 0 0 12 0.10 0.94 0.97 0.81 0.90 170 0.22 0.99 1.00
hilbert_transformer Xilinx 7 10 0 0 0 0 0 17 0.91 0.85 0.99 0.68 0.98 126 0.43 1.00 1.00
Lattice 7 10 0 0 0 0 0 17 0.89 0.90 0.99 0.74 0.96 126 0.46 1.00 1.00
@@ 5.1. Revisiting Research Questions
An overshadowing issue of the entire case study was the lack of adequate, openly available tooling to automatically investigate the recovered netlist at scale. Developing respective tooling, e.g., for netlist simulation, required immense efforts and was the main reason for many of the months spent on conducting the case study.
@@ 5.2. FPGA Threats & Defenses
A first line of defense is to impede the extraction of a gate-level netlist. Proprietary bitstream formats may initially raise the bar for a reverse engineering attack, but given increasing automation, this may be overcome within a reasonable time. This is inevitable since the user will always be able to translate HDL designs into bitstreams using EDA tools. As offered by many FPGA vendors, robust cryptographic schemes can ensure bitstream confidentiality, integrity, and authentication. By encrypting and signing a bitstream, attacks such as IP theft are impeded. However, such bitstream protections are not always provided on low-cost FPGAs and legacy devices, which continue to be used for decades.
@@
In light of recent attacks on bitstream encryption schemes (Ender et al., 2020; Moradi et al., 2011, 2012; Moradi and Schneider, 2016; Swierczynski et al., 2015b; Tajik et al., 2017), further precautions must be taken to harden the netlist against static and dynamic analysis by deteriorating those netlist properties that facilitate automated reverse engineering. Effective netlist obfuscation must aim at preventing the recovery of word-level structures as these structures are crucial for further analysis (Basiashvili et al., 2022). Dynamic reverse engineering is key to achieve an algorithmic understanding of larger circuits. Hence, to harden against such analysis techniques, non-simulatable primitives such as physical unclonable functions (Wendt and Potkonjak, 2014) or partial reconfiguration (Fyrbiak et al., 2018; Stolz et al., 2021; Fyrbiak et al., 2018) should be included in the design and merged into the control and data paths.
@@ 5.3. Related Work
Existing netlist reverse engineering techniques often only tackle isolated problems and have been evaluated on (sometimes outdated) open-source benchmarks rather than black-box designs. Hence, a comprehensive real-world case study is still lacking in the literature. For a summary of prior work, see Azriel et al. (Azriel et al., 2021).
@@ 5.4. Limitations and Future Work
Our case study, while focused on a single DSP design on a relatively small FPGA, still provides valuable insights. As we have shown in Section 4, our case study findings inform generalized techniques that are applicable to implementations across different architectures, designs, and sizes. Future research could build up on this work to explore more elaborate real-world designs on more complex FPGAs to further underline the threat potential of IP theft for FPGAs. Our work builds the foundation for such case studies by providing the techniques to dissect such a netlist at scale. Also, we have shown that there still is a need for better automated techniques for control logic extraction and analysis in gate-level netlists.
@@ Appendix B Details on the Maggie DSP
The input processing result and the closed-loop controller output are passed to a processing stage which implements a state observer, estimating the LRA’s position and velocity. In particular, Maggie appears to implement a proportional-integral (PI) observer (Beale and Shafai, 1989).
@@
The closed-loop controller takes a time series of desired actuator positions as its reference signal. Thus, as different haptic events are to be generated, this signal is changed. For instance, it is updated with the haptic feedback setting visible in Figure 12. Using the reference signal and the state observer outputs, the closed-loop controller accumulates the position and the velocity error (the desired velocity is obtained from numerical differentiation of the reference signal) and subtracts the integrated position estimation error. The particular combination of error components suggests that the control loop is a proportional-integral-derivative (PID)-controller (Ellis, 2002, 2012; Wu et al., 2017). Before being passed to the motor driver via I2S, a low-pass filter is applied to the output of the closed-loop controller.
@@ Appendix C Benchmarks
ibex (lowrisc, [n. d.]) Xilinx 3503 65 1937 0 1 727 6233
Lattice 6379 295 1963 16 0 78 8731
icicle (grahamedgecombe, [n. d.]) Xilinx 1525 65 1009 270 0 78 2947
Lattice 2705 262 983 20 0 59 4029
simple_risc_v (damdoy, [n. d.]) Xilinx 1617 76 1003 6 0 37 2739
Lattice 2876 274 1235 6 0 31 4422
canny_edge_detector (angelobacchini OpenCores, [n. d.]) Xilinx 2538 636 3713 24 0 72 6983
Lattice 3298 2170 4484 32 0 68 10052
fft64 (Sergiyenko and Uzenkov, [n. d.]) Xilinx 1722 212 1988 3 4 160 4089
Lattice 5362 2008 2556 14 0 124 10064
sha256 (secworks, [n. d.]) Xilinx 2196 104 1830 0 0 269 4399
Lattice 3935 346 2378 0 0 90 6749
hilbert_transformer (OpenCores, [n. d.]) Xilinx 304 68 587 0 0 85 1044
Lattice 445 225 569 2 0 78 1349

### v3

@@ 4.5. Guided Symbolic Execution
Starting at the cycle of interest, our guided SE approach first constructs the Boolean function describing the combinational sub-graph in front of a net. Next, it replaces all previously identified control signals with concrete values from the simulation while keeping all other variables symbolic. Each variable left in the Boolean function now corresponds to an output net of a sequential gate. If the sequential gate is not one of our endpoints, it finds the last cycle in which the gate was updated and continues back-tracing from the data inputs of the gate at that cycle. This process is repeated until the resulting Boolean function only contains variables corresponding to the output nets of our endpoints. To resolve potential recursive dependencies, our guided SE approach introduces intermediate variables to break up these dependencies when they are detected. To achieve this, we automatically identify sequential loops within the netlist and assign intermediate variables to the registers at the outputs of these loops.
@@
Applying guided SE is highly specific to the target implementation and the goals of the reverse engineer. Therefore, it is impossible to express its effectiveness in numbers. Still, we applied guided SE in a white-box case study and verified the results against a ground truth in Section 4.6 to demonstrate correctness. In a second evaluation step, we applied guided SE on the Maggie netlist to highlight its efficiency and scalability. To this end, we semi-automatically extracted a Python script describing the circuit’s algorithmic behavior during execution after initially performing the same task manually in Section 3.3. Thereby, we reduced the required effort from multiple weeks to merely a matter of days.
@@ 4.6. White-Box Case Study
To assess the applicability of our workflow, we conducted a white-box case study on a signal-processing design similar to Maggie. This additional study is conducted because it enables a comparison between the recovered design and a known ground truth, which is impossible in the black box setting. To this end, we synthesized an open-source Hilbert transformer design (OpenCores, [n. d.]) for a Xilinx 7-series FPGA and then extracted an algorithmic abstraction from the resulting gate-level netlist using only our generalized techniques.
@@
The initial netlist comprised 425 combinational and 619 sequential gates. After netlist pre-processing, we ended up with 1025 gates in total, 374 of them combinational and 651 sequential. Next, we recovered word-level structures using our generalized algorithms. The algorithm classified all 17 carry chains into 10 subtractions and 7 additions, identified 36 16-bit registers, and annotated 134 multi-bit pin groups to registers and arithmetic structures. When comparing against the ground truth, DANA achieved an NMI of 1.00, and our bit-order propagation assigned a correct bitorder to all pin groups. In total, we automatically assigned over 95% of all gates to word-level structures.
@@
ibex Lattice 5 0 4 0 0 0 0 9 0.05 0.98 0.98 0.97 0.96 168 0.27 0.40 1.00
Xilinx 1 0 2 0 0 0 3 6 0.01 0.98 0.98 0.96 0.97 148 0.16 0.89 1.00
icicle Lattice 2 0 7 0 0 1 0 10 0.12 0.86 0.92 0.89 0.87 112 0.31 0.62 0.94
Xilinx 1 1 4 0 0 0 1 7 0.05 0.91 0.94 0.88 0.95 346 0.79 0.87 1.00
simple_risc_v Lattice 3 1 4 0 0 6 0 14 0.13 0.99 0.98 0.98 0.96 139 0.33 0.95 1.00
Xilinx 3 1 4 0 0 2 0 10 0.07 0.98 0.98 0.99 0.99 117 0.29 0.91 1.00
canny_edge_detector Lattice 68 0 1 5 35 5 4 118 0.62 0.81 0.83 0.70 0.77 826 0.44 0.78 1.00
Xilinx 57 0 1 3 48 3 6 118 0.77 0.81 0.81 0.61 0.62 620 0.46 0.90 0.99
fft64 Lattice 93 4 7 0 4 0 1 109 0.40 0.90 0.92 0.81 0.87 604 0.59 0.94 0.96
Xilinx 25 5 0 0 4 0 6 40 0.27 0.88 0.95 0.70 0.90 351 0.37 0.97 0.97
sha256 Lattice 11 0 1 0 0 0 0 12 0.10 0.94 0.95 0.81 0.88 170 0.22 0.99 1.00
Xilinx 10 0 0 0 0 0 3 13 0.27 0.91 0.96 0.75 0.87 134 0.25 0.97 1.00
hilbert_transformer Lattice 7 10 0 0 0 0 0 17 0.89 0.90 0.98 0.74 0.96 126 0.46 1.00 1.00
Xilinx 7 10 0 0 0 0 0 17 0.91 0.85 0.99 0.68 0.98 126 0.43 1.00 1.00
@@ 5.1. Revisiting Research Questions
An overshadowing issue of the entire case study was the lack of adequate, openly available tooling to automatically investigate the recovered netlist at scale. Developing respective tooling, e.g., for netlist simulation, required immense effort and was the main reason for many of the months spent on conducting the case study.
@@ 5.2. FPGA Threats & Defenses
A first line of defense is to impede the extraction of a gate-level netlist. Proprietary bitstream formats may initially raise the bar for a reverse engineering attack, but given increasing automation, this may be overcome within a reasonable time. This is inevitable since the user will always be able to translate HDL designs into bitstreams using EDA tools. As offered by many FPGA vendors, robust cryptographic schemes can ensure bitstream confidentiality, integrity, and authentication. By encrypting and signing a bitstream, attacks such as IP theft are impeded. However, such bitstream protections are not always provided on low-cost FPGAs and legacy devices, which continue to be used for decades. For example, the Lattice iCE40 FPGA used in iPhone 7 does not feature any bitstream protections. Hence, Apple could not simply enable bitstream encryption to prevent reverse engineering attacks.
@@
In light of recent attacks on bitstream encryption schemes (Ender et al., 2020; Moradi et al., 2011a, 2012; Moradi and Schneider, 2016; Swierczynski et al., 2015b; Tajik et al., 2017), further precautions must be taken to harden the netlist against static and dynamic analysis by deteriorating those netlist properties that facilitate automated reverse engineering. Effective netlist obfuscation must aim at preventing the recovery of word-level structures as these structures are crucial for further analysis (Basiashvili et al., 2022). Dynamic reverse engineering is key to achieve an algorithmic understanding of larger circuits. Hence, to harden against such analysis techniques, non-simulatable primitives such as physical unclonable functions (Wendt and Potkonjak, 2014) or partial reconfiguration (Fyrbiak et al., 2018; Stolz et al., 2021; Fyrbiak et al., 2018) should be included in the design and merged into the control and data paths.
@@ 5.3. Related Work
Existing reverse engineering techniques often only tackle isolated problems (such as bitstream reverse engineering or word-level reconstruction) and have been evaluated on (sometimes outdated) open-source benchmarks rather than black-box designs. Hence, a comprehensive end-to-end real-world case study is still lacking in the literature.
FPGA Reverse Engineering
When reverse engineering an FPGA, the attacker commonly starts at the bitstream. A comprehensive overview of the availability of bitstream encryption and known attacks thereon is given in Appendix F. The process of reverse engineering the proprietary bitstream format is well understood (Ender et al., 2019; Ziener et al., 2006; Note and Rannaud, 2008; Benz et al., 2012; Ding et al., 2013; Zhang et al., 2023) and a variety of respective open-source tools exist (Alliance, 2023b; Kashani et al., 2022; Note, 2008; Alliance, 2023a; Pham et al., 2017). Ideally, such methods result in a database that allows the conversion of a bitstream into an error-free gate-level netlist describing the implementation on the analyzed FPGA. While we discuss existing netlist reverse engineering approaches below, none of the aforementioned works deals with the analysis of such an extracted gate-level netlist. However, numerous publications have discussed how to meaningfully manipulate the bitstream without reverse engineering the netlist. They often operate on the bitstream itself, i.e., without converting it to a netlist beforehand. For example, Swierczynski et al. (Swierczynski et al., 2015a) demonstrate targeted bitstream manipulations to weaken cryptographic primitives and mount an attack on a real-world device (Swierczynski et al., 2017). Similarly, Moraitis et al. (Moraitis and Dubrova, 2020) attack an FPGA implementation by replacing LUT configuration strings. In another case study, Kataria et al. (Kataria et al., 2019) defeat Cisco router security measures by manipulating the bits corresponding to the FPGA input/output (I/O). Swierczynski et al. (Swierczynski et al., 2018) and Engels et al. (Engels et al., 2023) automate such manipulations and demonstrated that they (under certain conditions) can be mounted even on encrypted bitstreams.
Netlist Reverse Engineering
For a comprehensive summary of prior work on netlist reverse engineering, see Azriel et al. (Azriel et al., 2021).
@@ 5.4. Limitations and Future Work
Our case study, while focused on a single DSP design on a relatively small FPGA, still provides valuable insights. As we have shown in Section 4, our case study findings inform generalized techniques that are applicable to implementations across different architectures, designs, and sizes. Future research could build up on this work to explore more elaborate real-world designs on more complex FPGAs to further underline the threat potential of IP theft for FPGAs. Due to the very limited amount of resources available on Lattice iCE40 FPGAs, the benchmarks used for evaluation are quite small in terms of logic gates. We are optimistic that our automated approaches scale well, even for larger netlists, while manual approaches naturally reach their limits rather quickly. Our work builds the foundation for such case studies by providing the techniques to dissect such a netlist at scale. Also, we have shown that there still is a need for better automated techniques for control logic extraction and analysis in gate-level netlists.
@@ Appendix B Details on the Maggie DSP
The input processing result and the closed-loop controller output are passed to a processing stage that implements a state observer, estimating the LRA’s position and velocity. In particular, Maggie appears to implement a proportional-integral (PI) observer (Beale and Shafai, 1989).
@@
The closed-loop controller takes a time series of desired actuator positions as its reference signal. This signal determines the haptic event to be generated and, therefore, is reconfigurable. For instance, it is updated with the haptic feedback setting visible in Figure 12. Using the reference signal and the state observer outputs, the closed-loop controller accumulates the position and the velocity error (the desired velocity is obtained from numerical differentiation of the reference signal) and subtracts the integrated position estimation error. The particular combination of error components suggests that the control loop is a proportional-integral-derivative (PID)-controller (Ellis, 2002, 2012; Wu et al., 2017). Before being passed to the motor driver via I2S, a low-pass filter is applied to the output of the closed-loop controller. We also obtained specifics of the implementation such as the concrete PID values, infinite impulse response (IIR) filter coefficients, and initialization values. However, to protect Apple’s IP, we do not disclose such details.
@@ Appendix C Benchmarks
ibex (lowrisc, [n. d.]) Lattice 6379 295 1963 16 0 78 8731
Xilinx 3503 65 1937 0 1 727 6233
icicle (grahamedgecombe, [n. d.]) Lattice 2705 262 983 20 0 59 4029
Xilinx 1525 65 1009 270 0 78 2947
simple_risc_v (damdoy, [n. d.]) Lattice 2876 274 1235 6 0 31 4422
Xilinx 1617 76 1003 6 0 37 2739
canny_edge_detector (angelobacchini OpenCores, [n. d.]) Lattice 3298 2170 4484 32 0 68 10052
Xilinx 2538 636 3713 24 0 72 6983
fft64 (Sergiyenko and Uzenkov, [n. d.]) Lattice 5362 2008 2556 14 0 124 10064
Xilinx 1722 212 1988 3 4 160 4089
sha256 (secworks, [n. d.]) Lattice 3935 346 2378 0 0 90 6749
Xilinx 2196 104 1830 0 0 269 4399
hilbert_transformer (OpenCores, [n. d.]) Lattice 445 225 569 2 0 78 1349
Xilinx 304 68 587 0 0 85 1044
@@ Appendix E White-Box Case Study Comparison
Appendix F Bitstream Encryption in the Wild
Here we provide an overview of the availability of bitstream encryption on many modern FPGAs from the top four FPGA vendors, namely AMD Xilinx, Intel Altera, Lattice Semiconductor, and Microchip. This list has been compiled from vendor documentation and to the best of our knowledge, but such resources may be incomplete or lack sufficient detail. Hence, we cannot guarantee correctness or completeness. Generally, older and cost-optimized FPGAs can be seen to lack bitstream encryption, while modern high-end FPGAs usually offer respective protections. While this list gives an intuition of how often vendors offer such features, we cannot give any numbers for how often these protections are actually enabled in the field. Historically, Xilinx FPGAs have received more attention from the academic community due to their market-leader position and availability to researchers, resulting in more successful attacks on their encryption schemes compared to other vendors. Hence, the lack of attacks on other vendors does not mean that their FPGAs are more secure when facing real-world attacks. All listed attacks strictly require access to the FPGA hardware.
Table 4. Overview of the availability of bitstream encryption across different vendors and FPGA families. We only state whether bitstream encryption is generally available for all (), some (), or none () of the FPGAs of a family, and whether attacks on the bitstream encryption have been demonstrated in academic literature. However, we do not evaluate the cryptographic strength of the encryption schemes.
Family Enc? Attacks Family Enc? Attacks
AMD Xilinx Virtex II (Moradi et al., 2011a) Spartan 6 (Moradi et al., 2011b; Moradi and Schneider, 2016)
Virtex 4 (Moradi et al., 2011b, 2012; Moradi and Schneider, 2016) Spartan 7 (Ender et al., 2020; Moradi and Schneider, 2016)
Virtex 5 (Moradi et al., 2011b, 2012; Moradi and Schneider, 2016) Artix 7 (Moradi and Schneider, 2016; Ender et al., 2020; Moradi and Schneider, 2016)
Virtex 6 (Ender et al., 2020; Moradi and Schneider, 2016) Kintex 7 (Moradi and Schneider, 2016; Ender et al., 2020; Tajik et al., 2017)
Virtex 7 (Ender et al., 2020; Moradi and Schneider, 2016) Ultrascale (all) (Ender et al., 2022; Lohrke et al., 2018)
Spartan 3 - Ultrascale+ (all) (Ender et al., 2022)
Intel Altera Agilex (all) - Stratix IV -
Arria (all) - Stratix V -
Max V - Stratix 10 -
Max 10 - Cyclone IV -
Stratix II (Moradi et al., 2013; Swierczynski et al., 2015b) Cyclone V -
Stratix III (Swierczynski et al., 2015b) Cyclone 10 -
Lattice iCE40 - CertusPro NX -
ECP2 - Avant -
ECP3 - Mach NX -
ECP5 - MachXO2 -
CrossLink - MachXO3 -
CrossLink NX - MachXO5 -
Certus NX - XP2 -
Microchip ProASIC - IGLOO 2 -
ProASIC 3 (Skorobogatov and Woods, 2012) Fusion -
PolarFire - SmartFusion -
IGLOO - SmartFusion 2 -